# Assessing Physics Students’ Scientific Argumentation using Natural Language Processing

### v1

@@
Scientific argumentation is an important science and engineering practice and a necessary 21st Century workforce skill. Due to the nature of large enrollment classes, it is difficult to individually assess students and provide feedback on their argumentation. The recent developments in Natural Language Processing (NLP) and Machine Learning (ML) may provide a solution. In this study we investigate methods using NLP and ML to assess and understand students argumentation. Specifically, we investigate the use of topic modeling to analyze student essays of argumentation after solving a problem in the recitation section of an introductory calculus-based physics course four semesters. We report on the emergent themes present in each semester.
@@
Winter Allen Department of Physics and Astronomy, Purdue University, 525 Northwestern Ave, West Lafayette, IN 47907, U.S.A.    Carina M. Rebello Dept. of Physics, Toronto Metropolitan University, 350 Victoria Street, Toronto, ON M5B 2K3, Canada    N. Sanjay Rebello Dept. of Physics and Astronomy / Dept. of Curriculum & Instruction, Purdue University, West Lafayette, IN 47907, U.S.A.
@@ 1 Introduction
For most students, their first official introduction to physics is in a high school classroom. This initial exposure often leaves a lasting impression, shaping how they view the subject for the rest of their lives. Many people walk away viewing physics as hard and not relevant to their day-to-day lives. However, learning physics encompasses far more than solving equations; it is about fostering a way of thinking that helps us understand the natural world. Oftentimes the disconnect between these aspects can be detrimental to student learning of physics. A core practice we want students to develop is the ability to solve problems; yet, the focus on plugging equations often overshadows this broader educational goal.
@@
Student problem solving is a relevant and rich topic in Physics Education Research (PER) and has been for many years [1, 2, 3]. Students often prioritize memorizing final answers over developing a deeper understanding of the problem-solving process [2, 4]. To foster growth in problem-solving, it is crucial not only to understand why students solve problems the way they do but also to help them reflect on their own problem-solving strategies, allowing them to develop as both learners and future scientists.
@@
Several research-based strategies have attempted to address the issues students have with problem solving [1]. We aim to address these fundamental issues through scientific argumentation. This idea is deeply rooted in philosophy [5] and has evolved significantly through educational research [5, 6]. Scientific argumentation is a proven strategy to help improve critical thinking that provides a schema for justifying the relevance of the retrieved knowledge in problem solving. To construct an argument students must justify their methods and decisions as they solved a problem, go through every step they took up to their solution, and provide evidence and reasoning for their process. In the context of problem-solving in physics, scientific argumentation involves not only an explanation of conceptual knowledge and methods but the ability to justify reasoning with empirical evidence and logical consistency. Within PER, scientific argumentation has been shown to enhance students’ ability to link theoretical knowledge with practical problem-solving skills [7]. This process encourages students to think critically about the methods they use and the evidence they gather, promoting skills that are essential for expert-like problem solving. Peer argumentation in physics classrooms also fosters collaborative learning, where students refine their ideas through group discussions and critique, further advancing their conceptual understanding and reasoning abilities [8]. The process of reflecting and evaluating one’s solution that the iterative nature of argumentation provides aligns with the goals of PER in promoting both content mastery and the development of scientific critical thinking.
@@
Scientific argumentation can be studied in both oral and written modalities. In this work, we focus on investigating scientific argumentation in students’ written essays in which they describe their strategies for solving problems. One of the major challenges that we face is systematically teaching and assessing student written essays in large enrollment classes due to the prohibitive time it takes for instructors or teaching assistants to read students’ written work and gauge their argumentation quality. Recent developments in Large Language Models (LLMs), Natural Language Processing (NLP), and Machine Learning (ML) may afford us the opportunity to address these challenges. ML is a subset of Artificial Intelligence (AI) that is the development of algorithms to detect patterns in datasets. NLP and LLM’s are subsets within ML that focus on and aid in the understanding of human language. Recently, there have been many studies on using LLMs, NLP, and ML in assessing student writing. [9, 10, 11, 12, 13]. Researchers have used these tools, such as unsupervised NLP, to assess student strategy essays [13] and explore the viability of utilizing LLMs in physics education [12, 11, 9, 10]. NLP and ML provide unique tools for analyzing large amounts of student text, allowing for the identification of patterns while still enabling qualitative, in-depth studies.
@@
To approach assessing students’ argumentation with ML in the context of problem-solving we have designed a study that introduces scientific argumentation through a series of scaffolds in the recitation portion of the course. We offer increasing levels of scaffolds through a four semester study. We then utilize unsupervised machine learning techniques to assess how student arguments change through the semesters.
2 Research Questions
The specific research questions (RQs) are:
RQ1: By introducing more levels of scaffolding throughout the semester, to what extent will students develop more complete, thoughtful arguments by the last module?
RQ2: To what extent can we use machine learning methods to assess student argumentation in the context of physics problem-solving?
@@ 3.1 Scientific Argumentation
We live in an age where misinformation is spread with a click of a button. We are constantly bombarded with dissenting information requiring strong critical thinking skills to wade through to the truth. Students often view science as another subject to memorize and regurgitate facts, rather than as a dynamic, evolving body of knowledge. This fundamental issue can be addressed by encouraging students to think critically about their work and engage in Scientific Argumentation (SA) in a classroom. It has been shown that by constructing explanations students may change their view of the nature of science [14]. As scientists, we know it is important to critically reflect on our work. Whether it be in research or solving a simple freshman level problem. We not only need to understand what decisions we made and why we made them in the problem-solving process, but we need to be able to present a well supported argument in support of our process. Research suggests that students tend to struggle with the idea of developing scientific arguments[15, 16], especially with finding appropriate evidence and constructing their reasoning [17, 18] and distinguishing between various elements of an argument [19, 20].
@@
Despite these difficulties, truly engaging in argumentation has shown to be beneficial to students. Not only does it aid in their own understanding of the problem, but it helps them learn to communicate and support their own findings [21]. To aid students in constructing arguments: argumentation scaffolds can elicit students’ participation in scientific argumentation [22] and a conducive learning environment can support students to solve problems, compare solutions, consider alternatives, and justify choices [17], [20]. Appropriate scaffolds include justification prompts [23] and question prompts [24]in instructional materials that help students articulate the rationale for their problem-solving steps and urge them to reason using evidence and justifications [25, 6] based on underlying principles [26]. However, most undergraduate physics courses do not facilitate scientific argumentation. Curricula that facilitate more expert-like problem solving can positively influence students’ epistemic beliefs and expectations around problem solving [27]. In more recent work, Rebello et al. [28, 7] found positive effects of using scientific argumentation in physics courses for future elementary teachers as well as future engineers.
@@
In recent years, scientific argumentation has been studied in various subjects: biology [29] [30], chemistry [31], and physics [32]. There are multiple ways to assess scientific argumentation [33] that rely on context in which the argument is presented, nature of the task being performed, and the specific learning goals [34]. For example, scientific argumentation in inquiry-based tasks may be assessed differently than in conceptual problem-solving activities. Moreover, these assessments must account for the complexity of scientific reasoning, the use of evidence, and how students articulate and justify their claims [35]. As a result, educators and researchers have developed diverse methodologies, such as Toulmin’s argumentation model [36, 5], which emphasize the structure of arguments, as well as rubrics that measure the quality of evidence and reasoning in students’ responses [37].
@@
Toulmin, a British philosopher, proposed breaking down arguments into six components: claim, grounds, warrant, qualifier, rebuttal, and backing.[5] A claim is the base purpose of an argument. The grounds are the evidence of the argument that support the claim. The warrant links the grounds to the claims. This can be explicitly stated or implied. The qualifier is used in wording the claim, while the rebuttal and backing are implied. The rebuttal acknowledges things that may contradict the claim. The backing establishes the relevance of the warrant. [36], [38] Toulmin’s model is based on arguments present primarily in law; however, this model has been particularly useful in examining how students construct arguments and justify their claims.
@@
McNeill and Krajcik [39] further adapted Toulmin’s model for use in science education by simplifying it into the Claim-Evidence-Reasoning (CER) framework. In their framework, the claim is an assertion or conclusion about a phenomenon, the evidence consists of scientific data supporting the claim, and the reasoning explains the relevance of the evidence. CER has become a popular framework in K-12 education, where students are encouraged to construct arguments using data to support their claims [40, 41]. Given the effectiveness of CER in K-12, there is a strong rationale for exploring its adaptation in undergraduate physics education, where developing students’ ability to argue scientifically can enhance their problem-solving and critical thinking skills.
@@
Scientific argumentation has been a rich topic to study in science education in K-12 classrooms in recent years. Erduran and Park [42] performed a search for manuscripts focusing on argumentation in secondary PER between 2003 and April 2022. During this period, they identified 13 published studies that explored scientific argumentation in secondary physics classrooms. These studies have provided critical insights into how students develop argumentation skills at the K-12 level, emphasizing the importance of scaffolding and structured support for helping students engage with complex scientific reasoning. However, despite the growing interest in this area, the relatively small number of studies over two decades suggests that research on argumentation, particularly in physics, is still in its early stages. Furthermore, these findings highlight the need for continued efforts to expand scientific argumentation research, particularly given the increasing emphasis on argumentation as a critical component of scientific literacy in modern curricula [37].
@@
Even less work has been done in undergraduate physics classrooms. While Erduran and Park [42] found 13 manuscripts on argumentation in secondary PER from 2003 to 2022, only nine manuscripts were published during the same period focusing on tertiary PER, highlighting the gap in research at the undergraduate level. In these nine manuscripts, there was a range of focus. The lack of study in this context does raise concerns as argumentation plays a critical role in developing advanced scientific reasoning and expert-like critical thinking skills, which are essential for physics students at the tertiary level as they transition into more complex problem-solving tasks that will aid them in the 21st century job market.
@@
In these nine manuscripts, a range of focus areas emerged. Some research concentrated on the epistemic tools learners use to construct scientific arguments [43], investigating how students draw upon their knowledge, reasoning, and evidence to build coherent, well-supported arguments. Other studies examined the relationship between student content knowledge and their performance on argumentation tasks [44], showing that students’ ability to engage in scientific argumentation often hinges on their depth of understanding of the underlying physical principles. Additionally, some research has explored the ways in which students use arguments across different disciplinary contexts, such as mathematics and physics, highlighting the challenges students face when applying argumentation strategies in varied subject areas [45].
@@
Despite these valuable insights, the limited volume of research on scientific argumentation in undergraduate physics classrooms points to the need for continued research in the topic. There is a lack of comprehensive studies that examine how scientific argumentation can be systematically integrated into traditional physics curricula. Most undergraduate physics courses tend to emphasize problem-solving over argumentation, leaving little room for students to engage in the kind of reflective, evidence-based reasoning that scientific argumentation requires. We aim to contribute by exploring how scientific argumentation can be effectively integrated into undergraduate physics recitations to improve students’ arguments.
@@ 3.2 Machine Learning
Machine learning is a branch of artificial intelligence that focuses on the development of statistical algorithms to understand patterns in unseen data. Two general sets of machine learning are unsupervised and supervised. Supervised machine learning relies on known, labeled data to train an algorithm to predict patterns in unseen data. Unsupervised machine learning focuses on discovering hidden patterns within datasets without explicit human guidance (training). Some popular examples of machine learning are clustering and topic modeling. In this paper, we chose to focus on unsupervised machine learning to analyze our data, due to to the large amount of student responses it would not be feasible to qualitatively extract emergent themes; nor would it be possible to label enough data to approach this project in a supervised or even semi-supervised way. We have worked with students for multiple years on argumentation, specifically on the same set of recitation problems, so we are confident in what we expect to see for the average student argument. Therefore, we feel confident in approaching this dataset with unsupervised techniques.
@@
Topic modeling is a technique that falls within the field of unsupervised machine learning. It is used to uncover hidden thematic structures in large collections of text. Rather than relying on labeled data, topic modeling algorithms infer topics based on patterns of word co-occurrence across documents.
One popular method of topic modeling is Non-Negative Matrix Factorization (NMF) [46], which is what we chose to use in this study. NMF is rooted in linear algebra and works by approximating an original document-term matrix as the product of two lower-dimensional matrices. The key constraint in NMF is that all values must be non-negative, which tends to produce more interpretable results—especially useful when analyzing human language, where negative values lack intuitive meaning.
Basically, suppose we have a document-term matrix $V$ of size $m\times n$, where $m$ is the number of documents and $n$ is the number of unique terms. NMF factorizes $V$ into two non-negative matrices: $W$ and $H$. $V$ represent the term document matrix, $W$ is an $m\times k$ matrix (where $k$ is the number of topics). Each row in $W$ represents a document’s distribution over topics. $H$ is a $k\times n$ matrix, where each row corresponds to a topic’s distribution over the terms. The idea is to find non-negative matrices W and H such that their product closely approximates the original matrix $V$. The rows of matrix $W$ give the document-to-topic weights, i.e., how much each document is associated with each topic. The rows of matrix $H$ give the topic-to-term weights, i.e., how much each term contributes to a given topic.
4 Methodology

### v2

@@
Scientific argumentation is a core science and engineering practice and a necessary 21st Century workforce skill. Due to the nature of large enrollment classes, it is difficult to individually assess students and provide feedback on their scientific argumentation. The recent developments in Natural Language Processing (NLP) and Machine Learning (ML) provide new opportunities to analyze large collections of student writing efficiently. In this study, we investigate how undergraduate students’ scientific argumentation evolves across four semesters of an introductory calculus-based physics course as increasingly structured argumentation scaffolds were introduced. We investigate the use of NLP and ML, specifically topic modeling, to analyze student scientific argumentation across those semesters. We report on the emergent themes present in each semester. Our findings show a clear shift in the thematic focus of student arguments corresponding to the level of scaffolding provided. In semesters with minimal scaffolding, students’ arguments emphasized procedural and surface-level features, while semesters with explicit scaffolds exhibited greater concentration around physics-principle-based themes. These results suggest that structured scaffolding supports students in constructing more conceptually grounded scientific arguments and highlights the potential of NLP and ML as scalable approaches for evaluate broad trends in students’ scientific argumentation.
@@
Winter Allen Dept. of Physics and Astronomy, Purdue University, 525 Northwestern Ave, West Lafayette, IN 47907, U.S.A.    Carina M. Rebello Dept. of Physics and Astronomy, Purdue University, West Lafayette, IN 47907, U.S.A.    N. Sanjay Rebello Dept. of Physics and Astronomy / Dept. of Curriculum & Instruction, Purdue University, West Lafayette, IN 47907, U.S.A.
@@ 1 Introduction
We live in an age where misinformation is spread with a click of a button. We are constantly bombarded with dissenting information requiring strong critical thinking skills to wade through to the truth. Students often view science as another subject to memorize and regurgitate facts, rather than as a dynamic, evolving body of knowledge. This fundamental issue can be addressed by encouraging students to think critically about their work and engage in scientific argumentation in a classroom.
@@
Student problem solving has been a relevant and rich topic in Physics Education Research (PER) for many years [doktor2014, Tulminaro2007, van1991learning]. Students often prioritize memorizing final answers over developing a deeper understanding of the problem-solving process [Tulminaro2007, dufresne1997solving]. To foster growth in problem-solving, it is crucial not only to understand why students solve problems the way they do but also to help them reflect on their own problem-solving strategies, allowing them to develop as both learners and future scientists.
@@
Several research-based strategies have attempted to address the difficulties that students have with problem solving [doktor2014]. Research has shown that scientific argumentation can improve problem solving [NUSSBAUM2003384]. This idea is deeply rooted in philosophy [Toulmin1958-TOUTUO-6] and has evolved significantly through educational research [Toulmin1958-TOUTUO-6, mcneill2010scientific]. Scientific argumentation is a proven strategy to help improve critical thinking that provides a schema for justifying the relevance of the retrieved knowledge in problem solving [asterhan2007effects, mcdonald2010influence, nussbaum2003argument]. To construct an argument students must justify their methods and decisions as they solved a problem, go through every step they took up to their solution, and provide evidence and reasoning for their process. In the context of problem-solving in physics, scientific argumentation involves not only an explanation of conceptual knowledge and methods but the ability to justify reasoning with empirical evidence and logical consistency. [cho2002effects]
Within PER, scientific argumentation has been shown to enhance students’ ability to link theoretical knowledge with practical problem-solving skills [rebello2019using]. This process encourages students to think critically about the methods they use and the evidence they gather, promoting skills that are essential for expert-like problem solving. Peer argumentation in physics classrooms also fosters collaborative learning, where students refine their ideas through group discussions and critique, further advancing their conceptual understanding and reasoning abilities [berland2011classroom]. The iterative nature of scientific argumentation i.e. the process of reflecting and evaluating one’s solution – aligns with the goals of promoting both content mastery and the development of scientific critical thinking.
@@
Scientific argumentation can be studied in both oral and written modalities [hakyolu2016interplay]. In this work, we focus on investigating scientific argumentation in written essays in which students describe their strategies for solving problems. One of the major challenges that we face is systematically teaching and assessing student written essays in large enrollment classes due to the prohibitive time it takes for instructors or teaching assistants to read students’ written work and gauge their argumentation quality. Recent developments in Large Language Models (LLMs), Natural Language Processing (NLP), and Machine Learning (ML) may afford us the opportunity to address these challenges. ML is a subset of Artificial Intelligence (AI) that is the development of algorithms to detect patterns in datasets. NLP and LLM’s are subsets within ML that focus on and aid in the understanding of human language. Recently, there have been many studies on using LLMs, NLP, and ML in assessing student writing [allen2025studentsperceptionslargelanguage, bralin2023analysis, sirnoorkar2025feedbackclicksintroductoryphysics, savage2025using]. [anand2023sciphyrag, wulff2024physics, polverini2023understanding, kieser2024using, tschisgale2023]. Researchers have used these tools, such as supervised and unsupervised NLP, to assess student strategy essays [munsell2021, tschisgale2023] and explore the viability of utilizing LLMs in physics education [kieser2024using, polverini2023understanding, anand2023sciphyrag, wulff2024physics]. NLP and ML provide unique tools for analyzing large amounts of student text, allowing for the identification of patterns while still enabling qualitative, in-depth studies.
We aim to address the gap in literature on longitudinal studies that investigate how students’ scientific argumentation, in the context of physics problem-solving, develops over time in response to the use of research-based instructional strategies to improve scientific argumentation. Due to the large datasets from large enrollment classes, it becomes increasingly difficult to track global student progress on argumentation in response to scaffolding. However, with the advent of ML, it is possible to leverage classic unsupervised machine learning (UML) techniques to study the longitudinal development of scientific argumentation skills for large numbers of students.
@@
To approach assessing students’ argumentation with ML in the context of problem-solving we have designed a study that introduces scientific argumentation through a series of scaffolds in the recitation portion of an introductory physics course. We offer increasing levels of scaffolds through a four semester study. We then utilize unsupervised machine learning techniques to assess how student arguments change through the semesters.
Our research question (RQ) is:
How does the quality of students’ scientific argumentation in the context of problem-solving in physics, evolve in response to scaffolding and fading over multiple weeks of instruction?
The current article is structured as follows: in the next section, we provide the background on prompt scientific argumentation and unsupervised machine learning. In Section III, we describe the study’s data, context, scaffolding, and analysis. In the following section, we present the results and discussion. Finally, in Section V, we then conclude with study’s implications, limitations, and directions for future research.
@@ 2.1 Scientific Argumentation
It has been shown that by constructing explanations students may change their view of the nature of science [bell2000scientific]. As scientists, we know it is important to critically reflect on our work, whether it be in research or solving a simple freshman level problem. We not only need to understand what decisions we made and why we made them in the problem-solving process, but we need to be able to present a well supported argument in support of our process. Research suggests that students tend to struggle with developing scientific arguments [berland2009making, kuhn1993science], especially with finding appropriate evidence and constructing their reasoning [berland2010learning, kuhn2010teaching] and distinguishing between various elements of an argument [forman1998you, jimenez2000doing].
@@
Despite these difficulties, engaging in scientific argumentation has shown to be beneficial to students. Not only does it aid in their own understanding of the problem, but it helps them learn to communicate and support their own findings [zohar2002fostering]. To aid students in constructing scientific arguments, argumentation scaffolds can elicit students’ participation in scientific argumentation [jonassen2010arguing] and a conducive learning environment can support students to solve problems, compare solutions, consider alternatives, and justify choices [berland2010learning], [jimenez2000doing]. Appropriate scaffolds include justification prompts [xun2004conceptual] and question prompts [cho2002effects]in instructional materials that help students articulate the rationale for their problem-solving steps and urge them to reason using evidence and justifications [christodoulou2014science, mcneill2010scientific] based on underlying principles [schworm2007learning]. However, most undergraduate physics courses do not facilitate scientific argumentation. Curricula that facilitate more expert-like problem solving can positively influence students’ epistemic beliefs and expectations around problem solving [wampler2013relationship]. In more recent work, Rebello et al. [rebello2019scaffolding, rebello2019using] found positive effects of using scientific argumentation in physics courses for future elementary teachers as well as future engineers.
@@
In recent years, scientific argumentation has been studied in various science disciplines including biology [anwar2019analyzing] [dorfner2018biology], chemistry [erduran2019argumentation], and physics [yang2023effects]. There are multiple ways to assess scientific argumentation [abi2011perceptions] that rely on context in which the argument is presented, nature of the task being performed, and the specific learning goals [sampson2008assessment]. For example, scientific argumentation in inquiry-based tasks may be assessed differently than in conceptual problem-solving activities. Moreover, these assessments must account for the complexity of scientific reasoning, the use of evidence, and how students articulate and justify their claims [crujeiras2017high]. As a result, educators and researchers have developed diverse methodologies, such as Toulmin’s argumentation model [toulmin2003uses, Toulmin1958-TOUTUO-6], which emphasize the structure of arguments, as well as rubrics that measure the quality of evidence and reasoning in students’ responses [berland2016epistemologies].
@@
Toulmin, a British philosopher, proposed breaking down arguments into six components: claim, grounds, warrant, qualifier, rebuttal, and backing. [Toulmin1958-TOUTUO-6] A claim is the base purpose of an argument. The grounds are the evidence of the argument that support the claim. The warrant links the grounds to the claims. This can be explicitly stated or implied. The qualifier is used in wording the claim, while the rebuttal and backing are implied. The rebuttal acknowledges things that may contradict the claim. The backing establishes the relevance of the warrant. [toulmin2003uses], [karbach1987using] Toulmin’s model is based on arguments present primarily in law; however, this model has been particularly useful in examining how students construct arguments and justify their claims.
@@
McNeill and Krajcik [mcneill2011supporting] further adapted Toulmin’s model for use in science education by simplifying it into the Claim-Evidence-Reasoning (CER) framework. In their framework, the claim is an assertion or conclusion about a phenomenon, the evidence consists of scientific data supporting the claim, and the reasoning explains the relevance of the evidence. CER has become a popular framework in K-12 education, where students are encouraged to construct arguments using data to support their claims [mcneill2008inquiry, wang2020scrutinising]. Given the effectiveness of CER in K-12, there is a strong rationale for exploring its adaptation in undergraduate physics education, where developing students’ ability to argue scientifically can enhance their problem-solving and critical thinking skills.
@@
Scientific argumentation has been a rich topic to study in science education in K-12 classrooms in recent years. Erduran and Park [erduran2023argumentation] performed a search for manuscripts focusing on argumentation in K-12 contexts between 2003 and April 2022. During this period, they identified 13 published studies that explored scientific argumentation in secondary physics classrooms. These studies have provided critical insights into how students develop argumentation skills at the K-12 level, emphasizing the importance of scaffolding and structured support for helping students engage with complex scientific reasoning. However, despite the growing interest in this area, the relatively small number of studies over two decades suggests that research on argumentation, particularly in physics, is still in its early stages. Furthermore, these findings highlight the need for continued efforts to expand scientific argumentation research, particularly given the increasing emphasis on argumentation as a critical component of scientific literacy in modern curricula [berland2016epistemologies].
@@
Even less research has been completed on scientific argumentation in undergraduate physics classrooms. While Erduran and Park [erduran2023argumentation] found 13 manuscripts on scientific argumentation in secondary settings from 2003 to 2022, only nine manuscripts were published during the same period focusing on tertiary settings, highlighting the gap in research at the undergraduate level. In these nine manuscripts, there was a range of focus. The lack of study in this context does raise concerns as argumentation plays a critical role in developing advanced scientific reasoning and expert-like critical thinking skills, which are essential for physics students at the tertiary level as they transition into more complex problem-solving tasks that will aid them in the 21st century job market.
@@
In these nine manuscripts, a range of focus areas emerged. Some research concentrated on the epistemic tools learners use to construct scientific arguments [cikmaz2021examining], investigating how students draw upon their knowledge, reasoning, and evidence to build coherent, well-supported arguments. Other studies examined the relationship between student content knowledge and their performance on argumentation tasks [jonsson2016student], showing that students’ ability to engage in scientific argumentation often hinges on their depth of understanding of the underlying physical principles. Additionally, some research has explored the ways in which students use arguments across different disciplinary contexts, such as mathematics and physics, highlighting the challenges students face when applying argumentation strategies in varied subject areas [van2019translating].
@@
Despite these valuable insights, the limited volume of research on scientific argumentation in undergraduate physics classrooms points to the need for continued research in the topic. There is a lack of comprehensive, longitudinal studies that examine how scientific argumentation can be systematically integrated into traditional physics curricula. Most undergraduate physics courses emphasize procedural problem-solving without incorporating active retrieval practices or argumentation-based activities that strengthen long-term memory and conceptual understanding. As a result, students have limited opportunities to engage in the kind of reflective, evidence-based reasoning that scientific argumentation requires. We aim to contribute to the existing literature by examining how scientific argumentation can be effectively integrated into undergraduate physics recitations to improve the quality of students’ written arguments. Further, given the large dataset of our investigation, we will employ systematic machine learning methods to track these changes across large textual datasets.
@@ 2.2 Machine Learning
Machine learning (ML) is a branch of artificial intelligence (AI) that focuses on the development of statistical algorithms to understand patterns in unseen data. ML includes both unsupervised and supervised [sarker2021machine] methods that are summarized below.
Supervised ML relies on known, labeled data to train an algorithm to predict patterns in unseen data. In contrast, unsupervised ML focuses on discovering hidden patterns within datasets without explicit human guidance [sarker2021machine]. Some popular examples of machine learning are clustering (i.e. K-Means [macqueen1967multivariate], DBSCAN [ester1996density], HDBSCAN [campello2013density] and topic modeling (i.e. LDA [blei2003latent], LSA [landauer1997solution], and NMF[lee1999learning]).
In this study, we used unsupervised machine learning to analyze our data. This decision is motivated by several factors. First, the large sizer of our data set makes it prohibitively time consuming to qualitatively extract emergent themes in a way that is reproducible and reliable. Second, although we have expectations of student arguments that we expect to see, based on the research-based strategies we have incorporated [rebello2019scaffolding, uruena2017impact, rebello2013transfer, rebello2013exploring]; we aim to capture all possible emergent themes at scale. Therefore, while supervised approaches could, in principle, be used to label and predict features of student arguments, doing so would require substantial manual annotation and would likely limit the analysis to predefined categories (e.g., binary classifications). Unsupervised techniques allow us to analyze large volumes of text data efficiently and discover meaningful structure without the extensive human labeling, thereby enabling a more exploratory and data-driven understanding of students’ written arguments.
@@
Topic modeling [lee1999learning, blei2003latent] is an unsupervised machine learning technique used to uncover hidden thematic structures within large collections of text. Rather than relying on labeled data, topic modeling algorithms infer topics based on patterns of word co-occurrence across documents [blei2012probabilistic].
One widely used method of topic modeling is Non-negative Matrix Factorization (NMF) [lee1999learning], which we employed in this study. NMF is rooted in linear algebra and works by approximating an original document-term matrix (e.g. student essay) as the product of two lower-dimensional matrices. The key constraint in NMF is that all values must be non-negative, which tends to produce more interpretable results—especially useful when analyzing human language, where negative values lack intuitive meaning.
A simple explanation of NMF is as follows: suppose we have a document-term matrix $V\in\mathbb{R}^{m\times n}_{+}$, where $m$ is the number of documents and $n$ is the number of unique terms. NMF factorizes $V$ into two non-negative matrices: $W\in\mathbb{R}^{m\times k}_{+}$ and $H\in\mathbb{R}^{k\times n}_{+}$ such that their product approximates the original matrix:
$V\approx WH$
Where $k$ represents the number of topics. Each row in $W$ represents a document’s distribution over topics. $H$ represents a topic’s distribution over the terms.
In summary:
$W_{m\times k}\times H_{k\times n}=V_{m\times n}$
where $m$ is the number of documents, $k$ is the number of topics, and $n$ is the number of unique terms.
3 Methods